# Leveraging AI-Generated Emotional Self-Voice to Nudge People towards their Ideal Selves

### v1

@@ 7.1. Summary of Findings
The intervention generally succeeded in inducing more positive affect across scenarios, partially supporting H1 and H3, and aligning with prior work showing positive affect for emotional synthetic voices (Cohn et al., 2020; Nass et al., 2001). However, the differences between ESV and other conditions were not as pronounced as initially hypothesized. This suggests that while ESV can be an effective tool, the act of reflection itself, regardless of modality, may have inherent benefits. Regarding resilience (H2), and confidence and motivation (H4), we observed improvements across conditions. Goal commitment (H5) showed a positive trend but didn’t reach statistical significance. This could be due to the short-term nature of our study, as commitment may require more time to solidify.
The contrasting nature of our two scenarios—past failure (Scenario 1) and future habit formation (Scenario 2)—revealed interesting differences in intervention effectiveness. Scenario 1 elicited more mixed emotional responses, with some participants experiencing negative feelings from revisiting past failures: “I just feel worse having spent time thinking about it” (P41), “Felt triggered a bit and somewhat unpleasant” (P21). This highlights the delicate nature of addressing past shortcomings and the need for careful framing in interventions. We also observed self-compassion as a strong predictor for resilience, which the ESV system could leverage. Scenario 2, focused on future habits, generally yielded more consistently positive responses. This forward-looking perspective seemed to align well with the ideal self concept, possibly because it allowed participants to envision improvement without the emotional baggage of past failures. These differences underscore the importance of context in behavioral interventions.
@@
Our technical characterization revealed that the generated text and voice content was consistently positive, especially for Scenario 2. The high ratings for voice familiarity, similarity, and quality in the ESV condition suggest that our voice cloning technology achieved its goal of creating a believable ”ideal self” voice. The parameterization findings offer intriguing insights into user preferences for idealized self-representations. The general preference for positivity and expressiveness, balanced with realism, suggests a sweet spot for effective interventions.
@@ 7.2. Implication for theory
The differential effects between past-focused and future-focused scenarios highlight the importance of framing in goal-related interventions, potentially extending our understanding of how temporal focus impacts goal pursuit.
@@ 7.3. Implications for designing generative self-voice systems for behavioral intervention
The differences between scenarios suggest that designers should consider whether past-focused or future-focused interventions are more appropriate for their specific use case. For example, the idea of a past failure brings out unpleasant feelings and conflates with ongoing goals. However, reframing past failure as a part of an ongoing goal might point the individual toward a more positive outlook. In general, designers can look into the psychology and behavioral science literature and test out different reflection prompts with the ESV system.
@@
Although the interaction presented in the prototypical system is very time-bounded (i.e., limited by what the study allows) and we cannot claim the longer-term effect on behavior change and habit formation, we believe that our system’s ability to generate emotional self-voice that can adapt to the individual’s scenario and ideal self can further adapt to their ever-changing needs in a longitudinal, real-world setting.
@@ 8. Ethical Considerations
This technology could pose privacy issues. However, this is not unique to this system. Any system that uses one’s image to generate hyper-realistic versions can pose an issue of impersonation and identity threat. Active research is being done on identifying deepfakes in the forms of image, audio, and video (Chen et al., 2020; Zhou and Lim, 2021; Rana et al., 2022; Groh et al., 2024).
While in this work we have shown the positive benefits of exposing one to a generated ideal-self, it is not hard to imagine how this technology can be used as a form of unpleasant marketing strategy or worse malicious manipulations that capitalize on one’s emotion and desire to better themselves. Specifically, self-voice implies the intervention is experientially personalized to an individual, which could make them more susceptible to manipulation. Thus, we cautioned the participants from blindly taking action based on the AI-generated responses, and we need to understand the long-term effect of using this intervention consistently.
@@ 9. Limitations & Future Work
Uncanny Valley of Generated Self Voice. Prior work has shown that an increase in voice familiarity can lead to eerie feelings due to uncanny valley (Chan et al., 2021; Holzman and Rousey, 1966; Kimura and Yotsumoto, 2018). Although not mentioned explicitly by the participants, we acknowledge the unpleasantness that comes with uncanny valley. Although the fact that people did not find the generated voices disturbing could either be the knowledge of the fact that it is an AI-generated version of their own voice (Hughes and Harrison, 2013).
Limitations of Self Voice Replication. The quality of the generation is affected by accents. We screened for people who speak fluent English, but this does not explicitly exclude accents. Although no participants explicitly mentioned the dissimilarity of the accent of the generated voice, in our early experiments, we did notice a non-American accent reduces the naturalness of the generated voice. Thus, the findings of this study might not generalize to cases where individuals’ accents are not well captured in the generated ESV.
@@
Range of Emotional Expressiveness. The current system has good emotional quality but the range of expression is limited to how the model was trained and the capabilities exposed to developers. The emotional range of the current implementation depends on the emotional TTS model’s interpretations of the emotion of the input text and the expressiveness of the samples provided for voice cloning. The range of emotional tonalities could be expanded and the preciseness of the expression, for example, being able to express different emotional attitude of the same text (Wu et al., 2024). Individuals can then experiment with hearing their own voice speaking the same text but in different attitudes.
Open-Source Systems. As mentioned in the previous paragraph, the models we used in the system are not “white-boxed”. Our proof-of-concept system prioritizes the speed and stability of the generation process. It would be better if the models were open-sourced. Our work contributes to the literature on the synthesis and interpretation of emotional speech, and we hope the findings will motivate the development of emotional speech generation for positive behavior change.
@@
Implication on Long-term Behavior Change. As an initial step, our study investigated how ESV affects one’s perception of an event, which can be a proxy of the direction of behavior change. Future work should investigate the efficacy of this technique on behavior change with a longitudinal study with actual behavioral tasks.
@@
Ideal-Self within Intrapersonal vs Interpersonal Contexts. We focused on scenarios in the self-development context. In reality, the concept of ideal self also applies to interpersonal communications (e.g., conflict resolution, asking for a raise). Future work should investigate the application of ESV in interpersonal contexts, where the ideal-self concept might adapt to others and the environment (Wille et al., 2018).
Ideal Self Voice vs Ideal Other Voice. We focused on the effect of the ideal self in this work. A natural extension of this work is to investigate the effect of an “ideal-other”, which is essentially a role model or someone else influential. Prior work has compared the effect of self-voice and other-voice, but the voice of an “ideal-other” adds another interesting dimension.
@@
Experiencing Other Types of Selves. In this work, we only explored the effect of experiencing the ideal self in the face of challenges. It would be interesting to explore the effect of experiencing other versions of the self, such as the current self or the “ought self”, which is who one or others think one should be (Higgins et al., 2013), which might be interesting in interpersonal contexts. Alternatively, exposing one to their current self might evoke negative emotions, but for some, negative emotions might drive more motivation and activation towards a goal. Future systems might have an adaptive strategy of motivating individuals depending on what they need in the moment.
Combining Other Modalities for Embodying Alternative Selves. Future work can explore the combination of emotional self voice with other modalities, such as seeing an ideal version of the self or embodying an ideal self in an immersive environment, to increase the overall vividness of an ideal self.
@@ 10. Conclusion
We introduce Emotional Self Voice (ESV), an innovative intervention designed to bridge the gap between one’s actual and ideal self. Our proof-of-concept system generates responses of an idealized self, delivered through a personalized, emotionally expressive cloned voice. Focusing on two critical aspects of goal achievement—overcoming failure and struggles with persistence—we demonstrate ESV’s potential to shift self-perception and behavior. Our study reveals that engaging with an ideal-self perspective generally increases positive affect and enhances key behavioral drivers such as resilience, confidence, motivation, and commitment across all conditions. While ESV didn’t consistently outperform other conditions, it demonstrated unique benefits, particularly in scenarios involving future habit formation. Participants showed high engagement with the ESV technology, expressing surprise and excitement at hearing their cloned voices articulate positive, motivational messages. The ability to fine-tune generated responses highlighted individual preferences for balancing positivity, expressiveness, and realism in self-talk. These findings underscore the potential of personalized, voice-based interventions in behavioral change contexts. As AI and voice cloning technologies continue to advance, ESV opens up exciting possibilities for creating adaptive, engaging interventions that support individuals in goal achievement and habit formation. Future research should explore the long-term effects of ESV, its applications across various behavioral domains, and the ethical implications of using AI-generated self-voices in psychological interventions. The significance of this work lies in its novel integration of advanced AI technologies with psychological theories of self-perception and behavior change, potentially revolutionizing how we approach personal development and mental health support. Our vision extends beyond the current implementation: we envision ESV as a ubiquitous, adaptive tool in everyday life, providing personalized support at critical moments of decision-making and self-reflection. Ultimately, we see ESV as a powerful tool for on-the-go reminders and interactive reflective processes, offering a new paradigm in personalized behavioral support that could significantly enhance individuals’ capacity for positive change and self-improvement.
@@ B.1. Demographics
B.2. Voice Cloning (ESV condition only

### v2

@@ 7.1. Summary of Findings
The intervention generally succeeded in inducing more positive affect across scenarios, partially supporting H1 and H3, and aligning with prior work showing positive effect for emotional synthetic voices (Cohn et al., 2020; Nass et al., 2001). However, the differences between ESV and other conditions were not as pronounced as initially hypothesized. This suggests that while ESV can be an effective tool, the act of reflection itself, regardless of modality, may have inherent benefits. Regarding resilience (H2), and confidence and motivation (H4), we observed improvements across conditions. Goal commitment (H5) showed a positive trend but didn’t reach statistical significance. This could be due to the short-term nature of our study, as commitment may require more time to solidify.
We compared the scenarios participants provided to the two prompts. While participants mentioned similar goals for both scenarios, such as health and fitness, the two sets of goals differ fundamentally in their scope and emotional landscape. The first set uncovers profound psychological struggles around significant life milestones, career advancement, and personal transformation, whereas the second set reveals tactical, routine-level obstacles centered on personal wellness. Essentially, the first represents high-stakes personal and professional aspirations, and the second represents incremental personal growth. Correspondingly, the contrasting nature of our two scenarios revealed interesting differences in intervention effectiveness. Scenario 1 elicited more mixed emotional responses, with some participants experiencing negative feelings from revisiting past failures: “I just feel worse having spent time thinking about it” (P41), “Felt triggered a bit and somewhat unpleasant” (P21). This highlights the delicate nature of addressing past shortcomings and the need for careful framing in interventions. We also observed self-compassion as a strong predictor for resilience, which the ESV system could leverage. Scenario 2, focused on future habits, more consistently yielded positive responses. This forward-looking perspective seemed to align well with the ideal-self concept, possibly because it allowed participants to envision improvement without the emotional baggage of past failures. These differences underscore the importance of context in behavioral interventions.
@@
Our technical characterization revealed that the generated text and voice content was consistently positive, especially for Scenario 2. The above-average ratings for voice familiarity, similarity, and quality in the ESV condition suggest that our voice cloning technology achieved its goal of creating a believable “ideal self” voice. The parameterization findings offer intriguing insights into user preferences for idealized self-representations. The general preference for positivity and expressiveness, balanced with realism, suggests a sweet spot for effective interventions.
@@ 7.2. Implication for theory
The differential effects between past-focused and future-focused scenarios highlight the importance of framing in goal-related interventions, potentially extending our understanding of how the temporal and cognitive focus impacts goal pursuit. For example, the idea of a past failure brings out unpleasant feelings and conflates with ongoing goals. However, reframing past failure as a part of an ongoing goal might point the individual toward a more positive outlook. In general, designers can look into the psychology and behavioral science literature and test out different reflection prompts with the ESV system.
@@ 7.3. Implications for designing generative self-voice systems for behavioral intervention
Future work should look into the effect of people’s prior experience with or attitudes towards AI and voice technologies. Prior work around AI chatbot therapist have found mixed results. Some showed positive attitude towards mental health chatbots (Abd-Alrazaq et al., 2021), while others found the attitudes varied by demographics(Benda et al., 2024). Compared to human-generated responses, AI-generated responses have been rated to have better perspective-taking and empathic concern (Yonatan-Leus and Brukner, 2024) while revealing nuanced differences in the focus and scope of the content (Li et al., 2024). Different from AI chatbot-based interventions, the role of AI in our system is not as salient given the responses were generated from a first-person perspective, and the awareness of AI as the source can alter the perception of the provided responses (Jain et al., 2024). We further see how the similar-attraction effect (Reeves and Nass, 1996) plays a role in hearing the generated response in the user’s own voice.
Although the interaction presented in the prototypical system is time-bounded (i.e., limited by what the study allows) and we cannot claim the longer-term effect on behavior change and habit formation, we believe that our system’s ability to generate emotional self-voice that can adapt to the individual’s scenario and ideal self can further adapt to their ever-changing needs in a longitudinal, real-world setting.
@@ 8. Ethical Considerations
This technology could pose privacy issues. However, this is not unique to this system. Any system that uses one’s information to generate hyper-realistic versions can pose an issue of impersonation and identity threat. Active research is being done on identifying deepfakes in the forms of image, audio, and video (Chen et al., 2020; Zhou and Lim, 2021; Rana et al., 2022; Groh et al., 2024).
In addition, the models we used in the system are not “white-boxed”. Our proof-of-concept system prioritizes the speed and stability of the generation process, and thus we opted for third-party, cloud-based models. It would be better if the models were open-sourced and locally hosted. The latter would also further improve the privacy of the system.
While in this work we have shown the positive benefits of exposing one to a generated ideal-self, it is not hard to imagine how this technology can be used as a form of unpleasant marketing strategy or worse, malicious manipulations that capitalize on one’s emotion and desire to better themselves. Specifically, self-voice implies the intervention is experientially personalized to an individual, which could make them more susceptible to manipulation. Thus, we cautioned the participants from blindly taking action based on the AI-generated responses, and we need to understand the long-term effect of using this intervention consistently.
@@ 9. Limitations & Future Work
Uncanny Valley of Generated Self Voice. Prior work has shown that an increase in voice familiarity can lead to eerie feelings due to uncanny valley (Chan et al., 2021; Holzman and Rousey, 1966; Kimura and Yotsumoto, 2018). Although not mentioned explicitly by the participants, we acknowledge the unpleasantness that comes with uncanny valley. The fact that people did not find the generated voices disturbing could be due to knowing that it was an AI-generated version of their own voice (Hughes and Harrison, 2013).
@@
Limitations of Self Voice Replication and Range of Emotional Expressiveness. The quality of the generation is affected by accents. We screened for people who speak fluent English, but this does not explicitly exclude accents. Although no participants explicitly mentioned the dissimilarity of the accent of the generated voice, in our early experiments, we did notice having a non-American accent reduces the naturalness of the generated voice. Thus, the findings of this study might not generalize to cases where individuals’ accents are not well captured in the generated ESV. The current system has good emotional quality but the range of expression is limited to how the model was trained and the capabilities exposed to developers. The emotional range of the current implementation depends on the emotional TTS model’s interpretations of the emotion of the input text and the expressiveness of the samples provided for voice cloning. The range of emotional tonalities could be expanded and the preciseness of the expression, for example, being able to express different emotional attitudes of the same text (Wu et al., 2024). Individuals can then experiment with hearing their own voice speaking the same text but with different attitudes.
@@
Implication on Long-term Behavior Change and Habituation. As an initial step, our study investigated how ESV affects one’s perception of an event, which can be a proxy of the direction of behavior change. Future work should investigate the efficacy of this technique on behavior change with a longitudinal study with actual behavioral tasks.
A follow-up study should explore the potential psychological impact of repeated exposure to one’s ideal self-voice, for example, changes in the user’s self-identity over time, whether the intervention promotes the internalization of ideal-self traits or cognitive dissonance.
In addition, we did not study the habituation effect towards the intervention, where users may become desensitized to hearing their own voice over time. While the study of habituation is beyond the scope of the study, we speculate the self-relevant stimuli, i.e., cloned self voice, have less habituation effect compared to other stimuli such as other voices (Rogers et al., 1977). On the other hand, the content might lead to habituation effect if the same cognitive reframing strategies are used repetitively. To mitigate this, the system could introduce tone, phrasing, or style variability to maintain engagement.
@@
Ideal-Self within Intrapersonal vs Interpersonal Contexts. We focused on scenarios in the self-development context. In reality, the concept of ideal self also applies to interpersonal communications (e.g., conflict resolution, asking for a raise). We tested the ESV on the interpersonal scenario for our pilot study, and we found that the generated responses of ESV varied more as the interpersonal challenges depended on the opponent’s personality and the relationship (positive or negative) between the user and the opponent. Future work should investigate the application of ESV in interpersonal contexts, where the ideal-self concept might adapt to others and the environment (Wille et al., 2018).
@@
Other Voice, Selves and Modalities. We focused on the effect of the ideal self in this work. A natural extension of this work is to investigate the effect of an “ideal-other”, which is essentially a role model or someone else influential. Prior work has compared the effect of self-voice and other-voice, but the voice of an “ideal-other” adds another interesting dimension. In this work, we only explored the effect of experiencing the ideal self. It would be interesting to explore the effect of experiencing other versions of the self, such as the current self or the “ought self”, which is who one or others think one should be (Higgins et al., 2013), which might be interesting in interpersonal contexts. Alternatively, exposing one to their current self might evoke negative emotions, but for some, negative emotions might drive more motivation and activation towards a goal. Future systems might have an adaptive strategy of motivating individuals depending on what they need in the moment. Future work can explore the combination of emotional self voice with other modalities, such as seeing an ideal version of the self or embodying an ideal self in an immersive environment, to increase the overall vividness of an ideal self.
@@ 10. Conclusion
We introduce Emotional Self Voice (ESV), a proof-of-concept system that generates responses of an idealized self, delivered through a personalized, emotionally expressive cloned voice. Focusing on two critical aspects of goal achievement—overcoming failure and struggles with persistence—we demonstrate ESV’s potential to shift self-perception and behavior. Our study reveals that engaging with an ideal-self perspective generally increases positive affect and enhances key behavioral drivers such as resilience, confidence, motivation, and commitment across all conditions. While ESV did not consistently outperform other conditions that only involved mental imagination or text-based interventions, it demonstrated unique benefits, particularly in scenarios involving future habit formation. Participants showed high engagement with the ESV technology, expressing surprise and excitement at hearing their cloned voices articulate positive, motivational messages. Future research should explore the long-term effects of ESV, its applications across various behavioral domains, and the ethical implications of using AI-generated self-voices in psychological interventions. The significance of this work lies in its novel integration of advanced AI technologies with psychological theories of self-perception and behavior change. We envision ESV as a ubiquitous, adaptive tool in everyday life, providing personalized support at critical moments of decision-making and self-reflection.
Acknowledgements.
The authors would like to thank A. Kapur and M. Cherep for their feedback on the manuscript, J. Shen for creating the teaser figure illustrations, and the study participants for their time.
@@ B.1. Demographics
B.2. Voice Cloning (ESV condition only)
@@ B.3. Pre-intervention
Alexithymia PSI audio PSI feeling Compassion Self discrepancy
Mean (SD) Mean (SD) Mean (SD) Mean (SD) Mean (SD)
Control 4.00 (1.33) 8.53 (1.47) 8.22 (1.55) 3.13 (0.54) 3.40 (1.63)
Text 3.99 (1.47) 8.83 (1.24) 8.30 (2.06) 3.07 (0.51) 3.85 (2.29)
ESV 4.31 (1.17) 8.75 (1.26) 7.53 (2.62) 2.80 (0.79) 5.30 (2.50)
Table 7. Demographics and Control Variables by Condition.
@@ B.4. Post-intervention
Appendix C Additional Study Results
C.1. Demographics by Condition
Table 7 shows the demographics breakdown for each condition.
C.2. Clustering Results
We performed a hierarchical clustering based on participants’ pre- and post-intervention responses to each scenario. Based on the clusters, we compare the control variables and demographic information of the sample population in each cluster to derive trends and insights. Below are the clusters for each scenario.
Figure 9. Pair plot visualizing the relationship between the outcomes across 3 clusters of scenario 1.
Figure 10. Demographic information of the clusters of Scenario 1.
Figure 11. Pair plot visualizing the relationship between the outcomes across 3 clusters of scenario 2.
Figure 12. Demographic information of the clusters of Scenario 2.